OpenStudy (anonymous):

I have done everything for this but #7. Please tell me what to do and please don't give me the answer thanks Consider the sequence 2,6,18,54... Let n = the term number in the sequence. Let A(n) = the value of the nth term of the sequence. 4. What is the common ratio of the sequence? 5. Complete each statement. a. A(1) = 2 = 2 * 3 = 2 * 3^? b. A(2) = 6 = 2 * 3 = 2 * 3^? c. A(3) = 18 = 2 * 3 * 3 = 2 * 3^? d. A(4) = 54 = 2 * 3 * 3 * 3 = 2 * 3^? 6. What is the relationship between the exponent of the base 3 and the value of n? 7. Complete the statement: A(n) = 2 * 3^?

OpenStudy (ash2326):

@connexus.com We need to complete this \[A(n)=2\times 3^{\large \text{^}}\] Ok, let's revisit the sequence 2, 6, 18, 54 Can you write the factors of each?

OpenStudy (anonymous):

I have done everything but #7 I don't under stand what it is asking for

OpenStudy (ash2326):

it's asking to complete the formula, Can you factor these? 2, 6, 18, 54...

OpenStudy (anonymous):

Yah its being multiplied by 3 each time but in other question it asked to do that so are you sure that's what its asking this is worth a big part of my grade so I need to get it all right

OpenStudy (ash2326):

yes, it's similar to 6. We just need to write the formula

OpenStudy (ash2326):

that will generate the terms as n is changed

OpenStudy (anonymous):

Its being multiplied by 3 each time so it would be 162,486,1458, & so on & so on

OpenStudy (ash2326):

that's right but we need to modify the formula suppose I had this 2, 4, 8, 16 I can write easily that \[A(n)=2\times 2^{n}\] where n=0, 1, 2, 3 and so on here also we need to similar

OpenStudy (anonymous):

Ok so for #7 I would write that A(n) = 2 * 3^n Were n = 0,1,2,3 and so on Is that what I need to put? If not I have no idea what to do lol

OpenStudy (ash2326):

yes, just A(n)

OpenStudy (anonymous):

Ok thanks

OpenStudy (ash2326):

yeah :)